OpenStudy (anonymous):

Hey guys is this C i def think it is pictures is in comment. i do give medals as well <3 <3 <3

OpenStudy (anonymous):

32 and 32 makes 64

TheSmartOne (thesmartone):

Sorry we can't help you on a quiz or a test. http://pages.citebite.com/c3n6a1q0a0mvl

TheSmartOne (thesmartone):

And fyi that answer is wrong and your work makes no sense...

OpenStudy (anonymous):

@Directrix

TheSmartOne (thesmartone):

What part confuses you?

TheSmartOne (thesmartone):

I'll solve a similar problem and see if you can solve this one y=x^2-121 For this it says y=0 So 0=x^2-121 x^2=121 x=\(\ \pm\) 11 Now can you solve that problem on your own?

OpenStudy (anonymous):

where did 11 come from?

TheSmartOne (thesmartone):

I am doing a similar problem.. not the same one. Cheating is not allowed..

OpenStudy (anonymous):

i know but i mean on ur prom where did u get 11

TheSmartOne (thesmartone):

Well I had 121 instead of 64...

OpenStudy (anonymous):

i know but u got 11 how?

TheSmartOne (thesmartone):

And \(\ \sqrt{121}= \pm 11\)

TheSmartOne (thesmartone):